### v3

# SMART: Submodular Data Mixture Strategy for Instruction Tuning

Abstract
Instruction Tuning involves finetuning a language model on a collection of instruction-formatted datasets in order to enhance the generalizability of the model to unseen tasks. Studies have shown the importance of balancing different task proportions during finetuning, but finding the right balance remains challenging. Unfortunately, there’s currently no systematic method beyond manual tuning or relying on practitioners’ intuition. In this paper, we introduce SMART (Submodular data Mixture strAtegy for instRuction Tuning) — a novel data mixture strategy which makes use of a submodular function to assign importance scores to tasks which are then used to determine the mixture weights. Given a fine-tuning budget, SMART redistributes the budget among tasks and selects non-redundant samples from each task. Experimental results demonstrate that SMART significantly outperforms traditional methods such as examples proportional mixing and equal mixing. Furthermore, SMART facilitates the creation of data mixtures based on a few representative subsets of tasks alone and through task pruning analysis, we reveal that in a limited budget setting, allocating budget among a subset of representative tasks yields superior performance compared to distributing the budget among all tasks. The code for reproducing our results is open-sourced at https://github.com/kowndinya-renduchintala/SMART.
SMART: Submodular Data Mixture Strategy for Instruction Tuning
H S V N S Kowndinya Renduchintala Media and Data Science Research Adobe Inc., India rharisrikowndinya333@gmail.com
Sumit Bhatia Media and Data Science Research Adobe Inc., India sumit.bhatia@adobe.com                        Ganesh Ramakrishnan Dept. of Computer Science & Engineering Indian Institute of Technology Bombay ganesh@cse.iitb.ac.in
1 Introduction
“Your ability to juggle many tasks will take you far.”
One of the main goals of artificial intelligence (AI) research is to build machines that can communicate (Turing, 1950), and an essential part of communication is to understand and follow instructions. Large Language Models (LLMs), which are pre-trained over massive text corpora on next-token-prediction objective, can perform a wide range of NLP tasks via “prompting” (Brown et al., 2020; Kojima et al., 2022; Almazrouei et al., 2023; Liu et al., 2023; Touvron et al., 2023).
Instruction Tuning (Wei et al., 2021; Sanh et al., 2021; Chung et al., 2022) is an approach that further enhances the instruction-following ability and generalizability of pre-trained LLMs to unseen tasks. It involves fine-tuning an LLM on a collection of instruction-formatted instances (encompassing multiple tasks) - each consisting of an instruction (or task description), an optional input, the corresponding output (the ground truth) and optionally a few demonstrations/examples. It is a special case of multitask learning where the LLM is finetuned on a collection of instruction-formatted multitask datasets Chung et al. (2022). Finetuning on multiple tasks simultaneously, allows the model to share and transfer information across tasks, resulting in a better common internal representation that is preferred by all tasks while suppressing task-dependent noise (Caruana, 1997). Consequently, the model learns to generalize to unseen tasks by discerning helpful cues from both implicitly and explicitly related tasks that it has previously seen.
The performance enhancement from instruction tuning is heavily contingent on data quality, data quantity, and task composition (Wang et al., 2023b). Studies by Iyer et al. (2022) and Longpre et al. (2023) have shown that while scaling the number of tasks is important, the relative proportion of various tasks (mixture weighting) merits as much attention for optimal instruction tuning. Intuitively, we want the model to see enough data for a given task that it can perform well on it, but not to see so much data that it memorizes the training set (Raffel et al., 2020).  Iyer et al. (2022) performed manual tuning of various benchmark proportions and decided on a final mixture, whereas Longpre et al. (2023) studied the impact of removing each benchmark from the finetuning mixture and relied on their practioners’ intuition from there on, to decide on the exact proportions of benchmarks. In this work, we would like to explore a more systematic approach to mixture weighting. Specifically, we are motivated by the fact that in a large multitask dataset like FLAN 2022 (Longpre et al., 2023), which has 1840 tasks, there will likely be many similar tasks leading to redundancies and not all of them may require sampling in equal proportions. For instance, there might be many tasks of the type Natural Language Inference (NLI), and it might be enough to sample relatively more instances from a few representative NLI tasks and less from the others. Furthermore, which samples we select from each task is also crucial because the samples should faithfully represent the task at hand. A random subset may fail to do this as it can miss out on essential corner cases.
With this context, we focus on the following two fundamental research questions (RQs) that form the basis for our subsequent inquiry:
  • (RQ1) Given a huge multitask instruction-tuning dataset and a limited fine-tuning budget which is defined by the total number of $(prompt,response)$ instances that can be used for fine-tuning, how do we divide this budget among thousands of tasks present in the dataset? i.e., how many instances to sample from each task? and which instances to sample from each task?
  • (RQ2) Can we go a step further and strategically prune some tasks altogether and only fine-tune on a small subset of representative tasks without hurting the performance? If yes, what is the nature of this subset?
To the best of our knowledge, there’s currently no principled approach to determining task compositions for instruction tuning, other than manual tuning and/or practioners’ intuition.
As a first step towards addressing both of the above RQs, we first define a common subset selection problem (more formally stated in Section 3) as follows - Given a huge collection of $M$ instruction-formatted task datasets, a task budget $M^{\prime}\leq M$ and a total budget ($N^{\prime}$) of $(prompt,response)$ pairs, which $M^{\prime}$ tasks to select? and how many instances to select from each of these $M^{\prime}$ tasks and which instances to select? Note that RQ1 is an instance of this problem where $M^{\prime}=M$.
Constrained Submodular Maximization (Section 2) proves to be a good model for discovering representative subsets (or coresets) of a massive training dataset (or ground set) that acts as surrogate (i.e., achieves similar performance) and are much better than uniformly-at-random subsets. Intuitively, this is because submodular functions model information in subsets, and hence maximizing a submodular function subject to a constraint yields non-redundant subsets of the ground set. An essential feature of this model is that it returns weighted subsets, i.e., each sample in the coreset comes with an associated score, which indicates how important the sample is.
Inspired by submodular functions, we propose our solution (Section 3) to the above subset selection problem for instruction tuning that works in two stages. In the first stage, we select a weighted subset of tasks from the full dataset where the weights will determine how many samples to select from each task. In the next stage, we select samples from each task based on the assigned task budgets. Note that the submodular functions used in each stage are not necessarily identical (Section 4.8).
The main contributions of our work can be summarized as follows:
  • We introduce SMART — a novel data mixture strategy for instruction tuning that models the data mixture problem (Section 3) as a sequence of two cardinality-constrained submodular maximization problems and offer empirical evidence that it outperforms both examples proportional and equal mixing baselines (Section 4) as well as the mixture weights proposed by Longpre et al. (2023).
  • Existing works like Longpre et al. (2023) have reported a continuous increase in performance upon increasing the number of tasks (though the gains themselves may be diminishing). However, we posit that this depends on the order in which new tasks are incorporated and show empirically that in the case of SMART mixtures, a performance peak is observed with an initial addition of few representative tasks and upon adding more and more tasks, the performance is not sustained (Section  4.6).
  • We find that the nature of instances that should be selected in each task (i.e, whether a representative or diverse subset) also depends on the total task budget, $M^{\prime}$ (Section 4.8). For higher $M^{\prime}$s, each task on average gets the relatively low budget and selecting representative samples is more important; however for lower $M^{\prime}$s, when there is sufficient enough budget for each task, the need for diversity dominates that of representation.
2 Background: Submodularity
Notations.
Let $f:2^{\mathcal{V}}\to\mathds{R}$ be a set function that assigns a value to every subset of the ground set $\mathcal{V}$. We use the notation $f(v|X)$ as a shorthand for $f(X\cup\{v\})-f(X)$ i.e., the incremental value gain of $v$ in the context of $X$.
Definition 1.
(Submodular Function) A given set function $f:2^{\mathcal{V}}\to\mathds{R}$ is submodular if for all $X,Y\subseteq\mathcal{V}$, where $X\subseteq Y$ and for all $v\notin Y$, the following inequality holds true:
$f(v|X)\geq f(v|Y)$
Intuitively, the definition states that — Adding an element to a smaller set $X$ yields more value gain than adding the same element to a superset $Y$ of $X$. Table 1 contains examples of three submodular functions — Facility Location (models representation), Log Determinant (models diversity), and Graph Cut (models a trade-off between representation and diversity controlled by the parameter $\lambda$).
Submodular Function $f(X)$
Facility Location $\underset{i\in\mathcal{V}}{\sum}{\underset{j\in X}{\max}{s_{ij}}}$
Graph Cut $\underset{i\in\mathcal{V},j\in X}{\sum}{s_{ij}}-\lambda\underset{i,j\in X}{\sum}{s_{ij}}$
Log Determinant $\log\det(\mathcal{S}_{X})$
Table 1: Examples of Submodular Functions. $\mathcal{V}$ is the ground set and $X\subseteq\mathcal{V}$. $s_{ij}$ is the similarity between two elements $i$ and $j$ of the ground set. $\mathcal{S}_{X}$ is the similarity matrix between the items in $X$. Facility Location models representation; Log Determinant models diversity and Graph Cut models a trade-off between representation and diversity (governed by the parameter $\lambda$).
Definition 2.
(Cardinality Constrained Submodular Maximization) Given a submodular function $f:2^{\mathcal{V}}\to\mathds{R}$ defined over the subsets of the ground set $\mathcal{V}$, the constrained submodular maximization problem involves finding $\mathcal{S}^{*}$ such that
$\mathcal{S}^{*}=\underset{\begin{subarray}{c}X\subseteq\mathcal{V}\\ |X|\leq N^{\prime}\end{subarray}}{\arg\max}{f(X)}$
The above cardinality-constrained submodular maximization problem is NP-complete Feige (1998). However, if $f$ is monotone submodular (i.e., $f(X)\leq f(Y)$ whenever $X\subseteq Y$), Nemhauser et al., 1978; Fisher et al., 1978 show that a simple greedy algorithm described in Algorithm 1 can be used to find an approximate solution $\mathcal{S}^{greedy}$, with a guarantee that $f(\mathcal{S}^{greedy})\geq(1-1/e)f(\mathcal{S}^{*})$. [^1] Assuming $P\neq NP$, this is the best approximation ratio that can be achieved by any polynomial time algorithm.
Algorithm 1 The Naïve Greedy
  Input: Ground Set ($\mathcal{V}$), Budget ($N^{\prime}$)
  $X_{0}\leftarrow\emptyset$;
  $\mathcal{S}\leftarrow[\,]$;
  $Gains\leftarrow[\,]$;
  for $i=0$ to ($N^{\prime}-1$do
     $e^{*}=\underset{v\in\mathcal{V}\setminus X_{i}}{\arg\max}{f(v|X_{i}})$;
     $g_{i+1}=f(e^{*}|X_{i})$;
     $X_{i+1}=X_{i}\cup\{e^{*}\}$;
     $\mathcal{S}\text{.append}(e^{*})$;
     $Gains\text{.append}(g_{i+1})$;
  end for
  return  $\mathcal{S},Gains$;
Remark.
The algorithm produces a weighted subset where the value gains themselves are the weights (i.e., $\underset{v\in\mathcal{V}\setminus X_{i}}{\max}{f(v|X_{i}})$ is the weight of $\underset{v\in\mathcal{V}\setminus X_{i}}{\arg\max}{f(v|X_{i}})$).
Algorithm 1 (also known as Naïve Greedy) requires $\mathcal{O}(N^{\prime}.|\mathcal{V}|)$ function evaluations which is costly in practice. Accelerated Greedy (Minoux, 2005), also known as Lazy Greedy, can be used instead, which leverages submodularity and offers a more efficient heap-based implementation of the same algorithm.
3 Approach
Inspired by submodularity (Section  2), in this section, we introduce a novel data mixture strategy, SMART, as a technique to solve the following subset selection problem introduced in Section  1:
Consider a collection $\mathcal{D}=\{\mathcal{T}_{1},\dots,\mathcal{T}_{M}\}$ of $M$ instruction-formatted task datasets where each $\mathcal{T}_{i}=\{(prompt_{ij},response_{ij})\}_{j=1}^{N_{\mathcal{T}_{i}}}$ consists of $N_{\mathcal{T}_{i}}$ $(prompt,response)$ pairs. Let $\sum_{i=1}^{M}N_{T_{i}}=N$. Given an $M^{\prime}\leq M$ and an $N^{\prime}\leq N$, how do we select a subset of tasks $\mathcal{D}^{\prime}=\{\mathcal{T}_{1}^{\prime},\dots,\mathcal{T}_{M^{\prime}}^{\prime}\}$(where $\mathcal{D}^{\prime}\subseteq\mathcal{D}$), and subsequently $\mathcal{S}=\{\mathcal{S}_{1},\dots,\mathcal{S}_{M^{\prime}}\}$ (where $\mathcal{S}_{j}\subseteq\mathcal{T}_{j}^{\prime}$) and $\sum_{j=1}^{M^{\prime}}|\mathcal{S}_{j}|=N^{\prime}$) such that efficiently fine-tuning on the subset $\mathcal{S}$ alone is (nearly) as effective as fine-tuning on the entire collection $\mathcal{D}$?
SMART models the above problem as a sequence of two cardinality-constrained submodular maximization problems. The first one is to select a weighted subset of $M^{\prime}$ tasks and the second one is to select instances - a total of $N^{\prime}$ instances from these tasks. The weights obtained in the first stage will be used to determine how many instances we sample from each task.
3.1 The Algorithm
We now give a detailed description of the two stages of SMART (summarized in Algorithm  2):
3.1.1 Stage-1: Weighted Task Subset Selection
In this first stage, given the instruction-tuning dataset $\mathcal{D}=\{\mathcal{T}_{1},\dots,\mathcal{T}_{M}\}$, our goal is to find $\mathcal{D}^{\prime}=\{\mathcal{T}_{1}^{\prime},\dots,\mathcal{T}_{M^{\prime}}^{\prime}\}$ where $\mathcal{D}^{\prime}\subseteq\mathcal{D}$, along with the instance budgets, $\{N_{1}^{\prime},\dots,N_{M^{\prime}}^{\prime}\}$, such that $\sum_{j=1}^{M^{\prime}}|N_{j}^{\prime}|=N^{\prime}$.
If $f_{1}$ is the submodular function that we use in this stage, $\mathcal{D}^{\prime}$ is given by:
$\mathcal{D}^{\prime}=\underset{\begin{subarray}{c}X\subseteq\mathcal{D}\\ |X|\leq M^{\prime}\end{subarray}}{\arg\max}{f_{1}(X)}$
To find the instance budgets ($N_{j}^{\prime}$s), we use the second-order Taylor-softmax operation (De Brebisson and Vincent, 2015) on the value gains obtained from the greedy algorithm, to compute a probability distribution which determines the probability with which instances will be sampled from a given task i.e., if $\{g_{1},\dots g_{M^{\prime}}\}$ are the value gains returned by the greedy algorithm, corresponding to the tasks $\{\mathcal{T}_{1}^{\prime},\dots,\mathcal{T}_{M^{\prime}}^{\prime}\}$, the instance budgets are given by
$N_{j}^{\prime}=\frac{(1+g_{j}+0.5g_{j}^{2})}{\sum_{k=1}^{M^{\prime}}(1+g_{k}+0.5g_{k}^{2})}\times N^{\prime}$
3.1.2 Stage-2: Instance Subset Selection
In this stage, given the subset of tasks, $\{\mathcal{T}_{1}^{\prime},\dots,\mathcal{T}_{M^{\prime}}^{\prime}\}$, and the instance budgets $\{N_{1}^{\prime},\dots N_{M^{\prime}}^{\prime}\}$ from the first stage, the goal is to actually select those many samples from each task. If $f_{2}$ is the submodular function used, the final subset $\mathcal{S}$ is given by
$\mathcal{S}=\bigcup_{j=1}^{M^{\prime}}\underset{\begin{subarray}{c}X_{j}\subseteq\mathcal{T}_{j}^{\prime}\\ |X_{j}|\leq N_{j}^{{}^{\prime}}\end{subarray}}{\arg\max}{f_{2}(X_{j})}$
Algorithm 2 The SMART Data Mixture Strategy
  Input: Task datasets $\mathcal{D}=\{\mathcal{T}_{1},\dots,\mathcal{T}_{M}\}$, Task Budget ($M^{\prime}$), Instance Budget ($N^{\prime}$), $f_{1}$, $f_{2}$# Each $\mathcal{T}_{i}=\{(prompt_{ij},response_{ij})\}_{j=1}^{N_{\mathcal{T}_{i}}}$
  $encoder\leftarrow\text{SentenceEncoder()}$()
  $prompt\_embeddings\leftarrow\text{dict()}$()
  $task\_embeddings\leftarrow\text{dict()}$()
  for $i=1$ to $M$ do
     $\mathbf{V}\leftarrow[\,]$
     for $j=1$ to $N_{\mathcal{T}_{i}}$ do
        $\mathbf{v}\leftarrow encoder\text{.encode}(prompt_{ij})$
        $\mathbf{V}\text{.append}(\mathbf{v})$
     end for
     $prompt\_embeddings[\mathcal{T}_{i}]=\mathbf{V}$
     $task\_embeddings[\mathcal{T}_{i}]=\dfrac{1}{N_{\mathcal{T}_{i}}}\sum\limits_{j=0}^{N_{\mathcal{T}_{i}}-1}\mathbf{V}[j]$
  end for
  $\mathcal{K}_{task}=\text{cos\_sim}(task\_embeddings)$
  $\mathcal{D}^{\prime},gains=\text{Greedy}(f_{1},\mathcal{D},\mathcal{K}_{task},M^{\prime})$# $\mathcal{D}^{\prime}=\{\mathcal{T}_{1}^{\prime},\dots\mathcal{T}_{M^{\prime}}^{\prime}\}$
  $probs\leftarrow\text{Taylor\_Softmax}(gains)$
  $\mathcal{S}^{\prime}\leftarrow[\,]$
  for $j=1$ to $M^{\prime}$ do
     $N_{j}^{\prime}=probs[j]\times N^{\prime}$
     $\mathcal{K}_{\mathcal{T}_{j}^{\prime}}=\text{cos\_sim}(prompt\_embeddings[\mathcal{T}_{j}^{\prime}])$
     $\mathcal{S}_{\mathcal{T}_{j}^{\prime}},gains=\text{Greedy}(f_{2},\mathcal{T}_{j}^{\prime},\mathcal{K}_{\mathcal{T}_{j}^{\prime}},N_{j}^{\prime})$
     $\mathcal{S}^{\prime}\text{.append}(\mathcal{S}_{\mathcal{T}_{j}^{\prime}})$
  end for
  $\mathcal{S}=\bigcup\limits_{j=1}^{M^{\prime}}\mathcal{S}^{\prime}[j]$
  return  $\mathcal{S}$
3.2 Obtaining Similarity Measures
The three submodular functions listed in Table 1 require computation of similarity measures ($s_{ij}$) between items in the ground set. So, all prompts in the dataset are first encoded using a sentence encoder. For instance subset selection (i.e., Stage-2), cosine similarity between prompt embeddings is used as the similarity measure and for weighted task subset selection (i.e., Stage-1), cosine similarity between task embeddings is computed, where the task embeddings are computed as the average prompt embeddings, following Vu et al. (2020). Although there are other sophisticated methods  (Achille et al., 2019; Zhou et al., 2022; Xi et al., 2023; Vu et al., 2021) for obtaining task embeddings, most of them depend on the LLM at hand.
3.3 Choosing $f_{1}$ and $f_{2}$
Each submodular function in Table 1 captures a different property: facility-location emphasizes representation, graph-cut balances representation and diversity, and log-determinant prioritizes diversity. We treat $f_{1}$ and $f_{2}$ as hyperparameters in our grid search in Section 4.8, exploring three options for each function. Determining the optimal $f_{1}$ and $f_{2}$ is part of our research question (RQ2). We discuss the findings of grid search and their qualitative implications for instruction tuning in Section 4.10.
4 Experiments
4.1 Finetuning Data
In all the experiments, the underlying ground set ($\mathcal{D}$) is FLAN 2022 (Longpre et al., 2023; Chung et al., 2022). The collection consists of the following five sub mixtures adding up to a total of 1840 tasks and 17,591,640 ($\sim$17.5M) instruction-formatted $(prompt,response)$ pairs:
  • FLAN 2021 (Wei et al., 2021)
  • T0  (Sanh et al., 2021)
  • NIV2 (Wang et al., 2022)
  • CoT (several chain-of-thought datasets)
  • Dialog (a few dialog datasets)
Each of the tasks comes in a variety of templates - zeroshot with and without answer options, fewshot with and without answer options.
SMART Data Mixture Creation
We encode prompts in the collection with GTE-large (Li et al., 2023b), a light-weight (340M parameters) BERT-based effective  (Muennighoff et al., 2022) sentence encoder for semantic textual similarity. Task embeddings are obtained by averaging the corresponding prompt embeddings (Section 3.2). We use SUBMODLIB (Kaushal et al., 2022), which has the necessary algorithms implemented, to obtain the weighted task subsets (Section  3.1.1) and the instances (Section  3.1.2). We distribute the instance budgets equally among task templates based on findings by Longpre et al. (2023) which showed that an equal number of zero-shot and few-shot templates yield the best performance on held-out tasks.
4.2 Finetuning Procedure
We evaluate SMART on three 7B parameter LLMs: Llama-2 Touvron et al. (2023), Falcon Almazrouei et al. (2023), and Mistral Jiang et al. (2023). We fine-tune each model for 1 epoch on a given data mixture with a learning rate of $52e-$ 5 for Llama-2 and Falcon, and $65e-$ 6 for Mistral. A batch size of $64$, weight decay of $0.1$, and cosine learning rate decay with a linear warmup for the initial $1\%$ training steps are employed. The experiments all ran on 8 NVIDIA A100-SXM4-80GB GPUs, utilizing Flash Attention Dao (2023) for memory efficiency and speeding up finetuning. Our code is open-sourced here.
4.3 Baselines
We mainly compare SMART with two baseline mixture strategies: Examples Proportional Mixture, Equal Mixture  (Raffel et al., 2020).
Examples Proportional Mixture (Baseline-1)
Instances are sampled in proportion to the size of each task’s dataset. This is equivalent to randomly sampling from the combined datasets.
Equal Mixture (Baseline-2)
Instances are sampled from each task with equal probability i.e., by dividing the total budget equally among tasks, and then uniformly sampling from each task.
We also compare SMART with mixture weights used by Longpre et al. (2023) but since their mixture weights apply to the five sub mixtures listed in Section 4.1 rather than individual tasks, we analyse this separately in Section 4.7.
$N^{\prime}$ Data Mix. MMLU-ZeroShot (Exact Match) BBH-Zeroshot (Exact Match) MMLU + BBH
STEM Humanities
Social
Sciences
Other
MMLU
FULL
NLP Algorithmic
BBH
FULL
(Weighted Avg.)
$25000$ EPM (Baseline-1) 30.82 46 46.71 43.05 40.63 40.59 25.55 31.67 38.05
EM (Baseline-2) 30.33 45.01 43.81 40.55 39.03 40.08 20.29 29.46 36.27
SMART (Ours) 32.22 50.41 50.14 46.85 43.73 38.85 24.16 30.05 39.8
$50000$ EPM (Baseline-1) 31.59 47.68 47.18 44.68 41.76 41.25 26.49 32.64 39.14
EM (Baseline-2) 35.22 49.58 51.01 48.13 44.99 41.96 22.83 31.24 41.04
SMART (Ours) 36.51 53.06 54.49 50.79 47.58 46.73 20.1 31.75 43.03
$100000$ EPM (Baseline-1) 32.66 50.6 51.37 47.18 44.25 43.38 26.36 33.48 41.16
EM (Baseline-2) 36.03 50.7 52.53 47.1 45.57 44.4 25.18 33.55 42.11
SMART (Ours) 37.36 55.38 55.47 52.95 49.11 47.26 24.22 34.66 44.96
$200000$ EPM (Baseline-1) 35.19 54.64 54.75 50.58 47.53 45.25 26.57 34.52 43.79
EM (Baseline-2) 38.6 54.05 54.72 51.47 48.68 41.36 24.75 32.28 43.96
SMART (Ours) 39.2 57.29 58.71 55.01 51.32 47.99 24.47 35.27 46.7
$400000$ EPM (Baseline-1) 38.16 56.53 56.99 52.56 49.85 48.72 26.04 36.49 46.01
EM (Baseline-2) 39.43 55.97 57.59 53.65 50.52 47.37 26.08 35.8 46.29
SMART (Ours) 39.77 57.39 60.17 54.79 51.77 49.25 26.35 37.43 47.65
17,591,640 Full FLAN 2022 42.44 59.1 61.82 55.1 53.43 50.7 27.6 38.11 49.03
Table 2: Comparison of SMART with baselines on MMLU-zeroshot and BBH-zeroshot for Llama-2-7b. EPM (Baseline-1) denotes Examples Proportional Mixture and EM (Baseline-2) denotes Equal Mixture. All the scores are Exact Matches. Weighted average on 57 MMLU tasks and 23 BBH tasks is reported in the last column. For baselines, exact matches are obtained by averaging across 3 fine-tuning runs.
4.4 Evaluation Protocol
We evaluate the fine-tuned models on two benchmark datasets: MMLU Hendrycks et al. (2020) with 57 tasks assessing world knowledge and problem-solving, and BBH Suzgun et al. (2022) with 23 challenging tasks from Big-Bench Srivastava et al. (2022). MMLU covers STEM, Humanities, Social Sciences, and Other (business, medical, and misc.) categories, while BBH includes both NLP and Algorithmic tasks [^2] The Algorithmic subcategory is so named because these tasks (e.g., 2-digit arithmetic) do not require an LLM to be solved.. Evaluation involves prompting the LLM directly and using Exact Match as the scoring metric. Responses are generated using the greedy decoding approach and they undergo basic post-processing steps (removing punctuation and lower-casing) before calculating exact match. For baseline data mixtures, 3 mixtures with different random seeds are created and the mean exact match of the 3 fine-tuning runs is reported.
4.5 Addressing RQ1 ($M^{\prime}=M$)
RQ1 is an instance of the subset selection problem defined in Section  3, where $M^{\prime}=M$, allowing all tasks for finetuning; however we still have a constraint ($N^{\prime}$) on total number of $(prompt,response)$ pairs. We use Graph-Cut and Facility Location functions (Table 1) in Stage-1 (Section 3.1.1) and Stage-2 (Section 3.1.2) of SMART respectively. This choice of $f_{1}$ and $f_{2}$ is determined via grid search discussed in Section 4.8. Table 2 contains comparison of SMART mixtures with baseline mixtures, on MMLU and BBH benchmarks, upon instruction fine-tuning Llama-2-7b on data mixtures generated by varying $N^{\prime}$ in $\{25000,50000,100000,200000,400000\}$. SMART data mixtures consistently perform better than both examples proportional mixtures and equal mixtures baseline.
Figure 1: Task-Scaling Curves in the case of Llama-2-7b for different $N^{\prime}$s, where X-axis is the number of tasks ($M^{\prime}$) and Y-axis is the weighted average of Exact Matches on MMLU and BBH.
4.6 Addressing RQ2 ($M^{\prime}<M$)
RQ2 is an instance of the subset selection problem defined in Section 3, where $M^{\prime}<M$ i.e., studying the effect of pruning some tasks altogether, on both SMART and the baseline data mixtures. We fine-tune Llama-2-7b on both the baseline and SMART data mixtures by varying the number of tasks ($M^{\prime}$) in $\{8,16,32,64,128,256,512,1024,1840\}$ and the total number of instances ($N^{\prime}$) in $\{25000,50000,100000,200000,400000\}$. Figure 1 depicts the task-scaling plots for each $N^{\prime}$. Baseline data mixtures’ performance steadily improves upon increasing the number of tasks, even though the gains are diminishing after a point. In contrast, SMART data mixtures yield optimal performance at an in-between point and this performance does not seem to sustain upon adding more and more tasks, suggesting that rather than increasing the tasks, focusing on representative tasks and sampling more instances from these might be more beneficial in low-budget scenarios. Even with an ample budget, scaling tasks should be done judiciously, as close to  97% of performance achievable by using the entire FLAN collection can be attained with just a subset of 16 representative tasks alone with $N^{\prime}=200000$.
$N^{\prime}$ FLANv2 Mix SMART
$25000$ 40.05 39.8
$50000$ 41.49 43.03
$100000$ 43.18 44.96
$200000$ 44.73 46.7
$400000$ 46.26 47.65
Table 3: Comparison of SMART with mixture weights of Longpre et al. (2023) for Llama-2-7b. The scores are weighted average of exact match on MMLU and BBH.
4.7 Comparison with FLANv2 Mix
We now compare SMART with FLANv2-mix i.e., by using the mixture weights suggested by Longpre et al. (2023) where the weights 40%, 32%, 20%, 5% and 3% are assigned to the 5 sub mixtures of FLAN 2022 — FLANv1, T0, NIV2, CoT and Dialog respectively and instances are randomly sampled from each sub mixture. Since this method only prescribes weights on sub mixtures and not on individual tasks, we only consider the case where $M^{\prime}=M$. Table 3 contains comparison of SMART vs FLANv2-mix for Llama-2-7b. FLANv2-mix seems to perform better than EPM (Baseline-1) and EM (Baseline-2) in some cases although SMART almost always outperforms FLANv2-mix.
4.8 Ablation Study: Submodular Function
$N^{\prime}=25000$ $N^{\prime}=50000$
$f_{1}$ $f_{2}$ FL GC LOGDET
FL 28.81 27.02 28.05
GC 39.8 40.84 40.61
LOGDET 38.83 35.52 41.38
$f_{1}$ $f_{2}$ FL GC LOGDET
FL 25.31 25.95 28.35
GC 43.03 40.98 42.28
LOGDET 41.97 40.67 41.84
$N^{\prime}=10000$ $N^{\prime}=200000$
$f_{1}$ $f_{2}$ FL GC LOGDET
FL 25.42 25.2 26.89
GC 44.96 43.92 43.61
LOGDET 43.63 42.25 42.81
$f_{1}$ $f_{2}$ FL GC LOGDET
FL 43.58 41.67 42.72
GC 46.7 45.91 46.21
LOGDET 43.75 43.86 43.82
Table 4: Grid Search on submodular functions $f_{1}$ and $f_{2}$ for Llama-2-7b where weighted average of exact match on MMLU and BBH are compared for different choices of $f_{1}$ and $f_{2}$. FL denotes Facility Location, GC denotes Graph Cut and LOGDET denotes Log-Determinant.
Figure 2: Varying $f_{2}$ when $f_{1}$ is Graph Cut
Both Stage-1 and Stage-2 of SMART require us to choose submodular functions $f_{1}$ and $f_{2}$ which are best treated as hyperparameters because of uncertainty with respect to which functions are best suited for instruction tuning (Section 3.3). We conduct a grid search on $f_{1}$ and $f_{2}$ using three functions from Table 1, by setting $M^{\prime}=M$ and varying $N^{\prime}$ in $\{25000,50000,100000,200000\}$. In this work, we use $\lambda=0.4$ for Graph-Cut. The grid search (summarized in Table 4) suggests that Graph Cut is optimal for task subset selection, while Facility Location is best for instance selection when $M^{\prime}=M$. The task scaling curves for each combination of $f_{1}$ and $f_{2}$ (for different $N^{\prime}$s) are present in Appendix B. However, in this section, we highlight the case where $f_{1}$ is Graph Cut and $f_{2}$ is varied. We find that optimal $f_{2}$ in this case also depends on $M^{\prime}$. For instance, in Figure 2 where $N^{\prime}=200000$, Facility Location performs the best at higher $M^{\prime}$, highlighting the importance of representation, while Graph Cut and Log Determinant show better performance at lower $M^{\prime}$, highlighting the importance of diversity. We hypothesize this is because — for higher $M^{\prime}$s, each task on average gets a relatively low budget and the need for representation dominates the need for diversity; however when there is sufficient enough budget for each task, i.e., for lower $M^{\prime}$s, the need for diversity takes over.
4.9 Ablation Study: LLM
We also test the effectiveness of SMART strategy on two other LLMs - Falcon-7B and Mistral-7B. Table 5 presents a comparison of SMART mixture with the baseline mixtures for Falcon and Mistral when $M^{\prime}=M$. For Falcon, we only report MMLU-zero-shot since Falcon gets an exact match of 0.0 for BBH-zero-shot even after fine-tuning. The task-scaling curves for these models are present in Appendix C.
$N^{\prime}$ Data Mix. Mistral Falcon
$25000$ EPM(Baseline-1) 52.69 23.09
EM(Baseline-2) 51.65 21.1
FLANv2-mix 53.24 26.16
SMART 53.33 22.73
$50000$ EPM(Baseline-1) 53.75 19.86
EM(Baseline-2) 53.16 22.96
FLANv2-mix 54.15 25.66
SMART 53.8 26.31
$100000$ EPM(Baseline-1) 54.24 20.88
EM(Baseline-2) 54.57 24.85
FLANv2-mix 54.14 27.47
SMART 53.75 27.22
$200000$ EPM(Baseline-1) 55.25 24.7
EM(Baseline-2) 54.71 24.39
FLANv2-mix 55.27 27.63
SMART 55.4 30.96
Table 5: Comparison of SMART with baselines for Mistral and Falcon. The scores correspond to weighted averages of exact match on MMLU and BBH.
4.10 Discussion
In Section 4.8, we saw that Graph Cut proves to be the most effective for selecting the weighted task subsets. Figures 678 of Appendix D contain t-SNE visualizations for task subsets selected by the three submodular functions listed in Table 1. Facility Location being a sum-max formulation, a single point is sufficient to represent a cluster and hence it gives more weightage to cluster centers and very less weightage to others. As a result, the submodular gains for first few selected points are very high and then the gains quickly become very small for other points in case of facility location. Log Determinant on the other hand predominantly selects diverse points not taking representation into account. Graph Cut, which models both, hence performs better than both Facility Location and Log Determinant for selecting tasks.
Further, in Figure 1 of Section 4.6 we saw that, at around 16 tasks, there is a peak in performance, after which there is a slight decline. To facilitate more investigation into what tasks are assigned more weightage, Table 6 and Table 7 list down the top-128 and last-128 tasks in the submodular ordering obtained using graph cut. While the former set contains more traditional NLP tasks like Natural Language Inference, Next Sentence Prediction, Question Answering, Summarization, etc ., the latter set mostly consists of tasks like Program Execution.
5 Related Work
5.1 Data for Instruction Tuning
Following Wang et al. (2023b), we summarize the related work in three categories — data quality, data quantity, and task composition.
Data Quantity
Research diverges on scaling instruction data quantity, with some advocating for limited data (Zhou et al., 2023; Chen et al., 2023a) to expose pretraining knowledge, while others argue for scaling up (Wei et al., 2021; Sanh et al., 2021). According to Ji et al. (2023); Dong et al. (2023); Yuan et al. (2023); Song et al. (2023), the impact of scaling varies across tasks and model abilities. AlShikh et al. (2023) also introduce a metric for instruction following ability, and suggest an early stopping criterion for instruction-tuning.
Data Quality
High-quality data is crucial in instruction tuning (Chia et al., 2023; Ding et al., 2023; Zhou et al., 2023). Wang et al. (2023a) use perplexity to select suitable instructions generated by models, while Li et al. (2023a) employ the language model itself to augment and curate high-quality training examples to improve its own performance. Cao et al. (2023) propose InstructionMining, utilizing natural language indicators to predict inference loss as a proxy for data quality without human intervention and select the best subset based on this. Chen et al. (2023b) introduce AlpaGasus, which uses a strong LLM to select high-quality subsets. Lu et al. (2023) and Madaan et al. (2024) also leverage the power of fine-tuned LLM itself to evaluate the quality of instructions. Attendu and Corbeil (2023) employs dynamic data subset selection by filtering out unimportant samples during finetuning, based on an extended EL2N metric (Paul et al., 2021; Fayyaz et al., 2022). Taori et al. (2023) propose #InsTag to assess instruction diversity in SFT datasets using ChatGPT. Additionally, Wan et al. (2023) propose Explore-Instruct, utilizing LLMs to actively explore domain-specific spaces and gather diverse instruction-tuning data. Wu et al. (2023) select new data points that are distinct from existing ones in the model embedding space, augmenting the training dataset iteratively to enhance diversity within subsets.
Task Composition
Many previous works show the benefit of scaling the number of tasks (Wei et al., 2021; Chung et al., 2022; Wang et al., 2022; Sanh et al., 2021). However, works like Iyer et al. (2022); Longpre et al. (2023) have also acknowledged that task balancing is also very important for effective instruction tuning. Dong et al. (2023) explore data composition across GSM8k, Code Alpaca, and ShareGPT datasets, finding differential impacts of data scaling on performance across abilities and propose a Dual-stage mixed fine-tuning strategy as a promising solution to activate multiple abilities efficiently. Ivison et al. (2022) identifies relevant multitask subsets based on the similarity between the pre-trained model’s representations, using a small amount of target task data. Yin et al. (2023) uses instruction representations for task selection, acting as a replay strategy to mitigate catastrophic forgetting and improve generalization in continual learning. Yue et al. (2023) construct math generalist models via instruction tuning on hybrid of chain-of-thought and program-of-thought rationales in math. Lou et al. (2023); Zhang et al. (2023) provide a survey of instruction tuning in general and Wang et al. (2023b) provide a detailed survey of data management for instruction tuning.
5.2 Submodularity for Subset Selection
Submodularity (Fujishige, 2005) has a long history in combinatorial optimization, game theory, economics etc (Edmonds, 1970; Lovász, 1983; Carter, 2001; Topkis, 1998). It has recently gained traction in machine learning where it has been used for data subset selection for machine translation (Kirchhoff and Bilmes, 2014), speech recognition (Wei et al., 2014; Mittal et al., 2022; Kothawade et al., 2023), efficient pre-training of language models (Renduchintala et al., 2023), active learning (Wei et al., 2015; Kothawade et al., 2021), hyperparameter tuning (Killamsetty et al., 2022), domain adaptation (Karanam et al., 2022), computer vision (Kaushal et al., 2019), continual learning (Tiwari et al., 2022) etc. For a more detailed review of submodularity and its applications, please refer to the survey by Bilmes (2022).
6 Conclusion & Future Work
In this paper, we introduced SMART — a novel data mixture strategy for instruction tuning that utilizes a submodular function to assign importance scores to tasks, determine the mixture weights, and also select non-redundant samples from each task. Further, we also reveal that in a low-budget setting, splitting the budget among a small subset of representative tasks yields superior performance when compared to dividing it among all tasks, which suggests that task scaling should be done more judiciously. Future work could explore making this method more model-specific (e.g., modify task embedding computation) and also possibly modify the approach for targeted instruction-data selection for creating expert LLMs that specialize in specific skills such as math, code etc.
7 Limitations
While the approach has its own advantages of computing the data mixture only once for a given dataset and using it for as many LLMs as one may wish, the approach might benefit from taking inputs from the language model as well and perform model specific instruction tuning. Secondly, SMART shows that there is an optimal number of tasks at which there is peak in performance observed. However, it doesn’t say anything on how to find the optimal point as it may depend on language model, the total budget and most importantly the underlying dataset.
8 Ethical Considerations
While instruction tuning leverages pretrained language models and encompasses similar considerations, we anticipate that this approach will predominantly yield positive outcomes. It offers an enhanced methodology for task balancing, potentially allowing for more cost-effective fine-tuning of large language models compared to conventional data mixture strategies.
Acknowledgements
The authors acknowledge the use of ChatGPT [^3] https://chat.openai.com/ for solely paraphrasing and summarizing some parts of the paper and declare that none of the generated content is presented in the paper without rigorous manual checking.
APPENDIX
Appendix A Code and Data
Our code we used for instruction tuning and for creating data mixtures is open-sourced at https://github.com/kowndinya-renduchintala/SMART. The code development utilized open-source tools, primarily relying on the HuggingFace library for model training with PyTorch as the underlying framework. Both PyTorch and HuggingFace are licensed under permissive licenses, with PyTorch under the BSD license and HuggingFace under the Apache 2.0 license. Additionally, submodular optimization was performed using SUBMODLIB, which is an openly accessible library on GitHub at https://github.com/decile-team/submodlib under the MIT license.
Appendix B Task Scaling Curves: Varying $f_{1}$ and $f_{2}$
Figure 3: Task Scaling Curves for various $f_{1}$ and $f_{2}$ combinations for Llama-2-7b
Figure 3 consists of task scaling curves for the 9 possible combinations (Section 4.8) of $f_{1}$ and $f_{2}$.
Appendix C Task Scaling Curves: Mistral, Falcon
Figure 4: Task Scaling Curves for Mistral-7B
Figure 5: Task Scaling Curves for falcon-7B
Figure 4 and Figure 5 consist of task scaling curves for Mistral-7B and Falcon-7B respectively (Section 4.9).
Appendix D Visualization of Task Subsets
Figure 6: t-SNE visualizations for task subsets of various sizes ($M^{\prime}$) selected by Facility Location
Figure 7: t-SNE visualizations for task subsets of various sizes ($M^{\prime}$) selected by Graph Cut
Figure 8: t-SNE visualizations for task subsets of various sizes ($M^{\prime}$) selected by Log Determinant
Rank Task
1 anli/r3:0.1.0
2 hellaswag:1.1.0
3 task1392_superglue_multirc_answer_verification
4 task955_wiki_auto_style_transfer
5 task381_boolq_question_generation
6 task1291_multi_news_summarization
7 task1295_adversarial_qa_question_answering
8 task519_aquamuse_question_generation
9 anli/r2:0.1.0
10 anli/r1:0.1.0
11 race_high_Select_the_best_answer_no_instructions_
12 cot_ecqa_ii
13 super_glue/rte:1.0.2
14 task1660_super_glue_question_generation
15 task339_record_answer_generation
16 task302_record_classification
17 stream_qed_ii
18 task870_msmarco_answer_generation
19 paws_wiki:1.1.0
20 super_glue/multirc:1.0.2
21 task603_wikitext-103_fill_in_the_blank
22 task887_quail_answer_generation
23 stream_qed
24 task380_boolq_yes_no_question
25 coqa:1.0.0
26 task1412_web_questions_question_answering
27 super_glue/cb:1.0.2
28 task1293_kilt_tasks_hotpotqa_question_answering
29 quail_context_question_description_answer_text
30 fix_punct
31 true_case
32 winogrande:1.1.0
33 glue/wnli:2.0.0
34 super_glue/record:1.0.2
35 cot_ecqa
36 quail_context_question_description_answer_id
37 task919_coqa_incorrect_answer_generation
38 task520_aquamuse_answer_given_in_passage
39 task1290_xsum_summarization
40 task1609_xquad_en_question_generation
41 squad/v1.1:3.0.0
42 task231_iirc_link_classification
43 task349_squad2.0_answerable_unanswerable_question_classification
44 wiki_dialog_ii
45 task1661_super_glue_classification
46 quail_context_description_question_text
47 adversarial_qa_droberta_answer_the_following_q
48 task644_refresd_translation
49 bool_q:1.0.0
50 task470_mrqa_question_generation
51 race_middle_Select_the_best_answer_no_instructions_
52 task770_pawsx_english_text_modification
53 adversarial_qa_dbidaf_answer_the_following_q
54 glue/qnli:2.0.0
55 task1558_jfleg_incorrect_answer_generation
56 task344_hybridqa_answer_generation
57 quoref_Guess_Title_For_Context
58 super_glue/copa:1.0.2
59 word_segment
60 task1340_msr_text_compression_compression
61 wiki_dialog
62 squad/v2.0:3.0.0
63 task1530_scitail1.1_sentence_generation
64 task225_english_language_answer_generation
Rank Task
65 task303_record_incorrect_answer_generation
66 gem/web_nlg_en:1.1.0
67 task768_qed_text_span_selection
68 adversarial_qa_dbert_answer_the_following_q
69 task1389_hellaswag_completion
70 task1294_wiki_qa_answer_verification
71 wiki_qa_Topic_Prediction_Answer_Only
72 task596_mocha_question_generation
73 task871_msmarco_question_generation
74 task1564_triviaqa_answer_generation
75 task1344_glue_entailment_classification
76 cosmos_qa:1.0.0
77 task1555_scitail_answer_generation
78 task1557_jfleg_answer_generation
79 gem/common_gen:1.1.0
80 task238_iirc_answer_from_passage_answer_generation
81 task233_iirc_link_exists_classification
82 super_glue/wic:1.0.2
83 task1345_glue_qqp_question_paraprashing
84 glue/qqp:2.0.0
85 glue/stsb:2.0.0
86 task595_mocha_answer_generation
87 task460_qasper_answer_generation
88 glue/mnli:2.0.0
89 task051_multirc_correct_answer_single_sentence
90 adversarial_qa_droberta_tell_what_it_is
91 task547_alt_translation_entk_en
92 quail_no_prompt_id
93 task311_race_question_generation
94 cot_sensemaking_ii
95 gem/dart:1.1.0
96 wiki_qa_Jeopardy_style
97 adversarial_qa_dbidaf_tell_what_it_is
98 task1593_yahoo_answers_topics_classification
99 quail_context_question_answer_description_text
100 cot_creak_ii
101 definite_pronoun_resolution:1.1.0
102 gigaword:1.2.0
103 super_glue/wsc.fixed:1.0.2
104 task234_iirc_passage_line_answer_generation
105 task556_alt_translation_en_ja
106 task604_flores_translation_entosn
107 adversarial_qa_dbert_tell_what_it_is
108 task310_race_classification
109 task1594_yahoo_answers_topics_question_generation
110 story_cloze/2016:1.0.0
111 task933_wiki_auto_style_transfer
112 gem/wiki_lingua_english_en:1.1.0
113 task1382_quarel_write_correct_answer
114 task1296_wiki_hop_question_answering
115 quail_context_description_question_answer_text
116 glue/cola:2.0.0
117 cot_strategyqa_ii
118 task1553_cnn_dailymail_summarization
119 openbookqa:0.1.0
120 task054_multirc_write_correct_answer
121 task1218_ted_translation_en_ja
122 quail_no_prompt_text
123 quail_context_question_description_text
124 task550_discofuse_sentence_generation
125 quail_context_question_answer_description_id
126 task1608_xquad_en_answer_generation
127 task1520_qa_srl_answer_generation
128 cos_e_v1.11_generate_explanation_given_text
Table 6: List of 128 most representative tasks in FLAN-2022 collection as ordered by the Graph Cut
Rank Task
1713 task261_spl_translation_es_en
1714 task1384_deal_or_no_dialog_classification
1715 task854_hippocorpus_classification
1716 task110_logic2text_sentence_generation
1717 task360_spolin_yesand_response_generation
1718 task148_afs_argument_quality_gay_marriage
1719 task499_extract_and_add_all_numbers_from_list
1720 task176_break_decompose_questions
1721 task085_unnatural_addsub_arithmetic
1722 task108_contextualabusedetection_classification
1723 task472_haspart_classification
1724 task856_conv_ai_2_classification
1725 task600_find_the_longest_common_substring_in_two_strings
1726 task150_afs_argument_quality_gun_control
1727 task1508_wordnet_antonyms
1728 task183_rhyme_generation
1729 task488_extract_all_alphabetical_elements_from_list_in_order
1730 task682_online_privacy_policy_text_classification
1731 task1425_country_iso_numeric
1732 task756_find_longert_substring_and_return_all_unique_alphabets_in_it
1733 task1585_root09_hypernym_generation
1734 task958_e2e_nlg_text_generation_parse
1735 task584_udeps_eng_fine_pos_tagging
1736 task1319_country_by_barcode_prefix
1737 task1507_boolean_temporal_reasoning
1738 task509_collate_of_all_alphabetical_and_numerical_elements_in_list_separately
1739 task064_all_elements_except_first_i
1740 task130_scan_structured_text_generation_command_action_long
1741 task365_synthetic_remove_vowels
1742 task149_afs_argument_quality_death_penalty
1743 task210_logic2text_structured_text_generation
1744 task1495_adverse_drug_event_classification
1745 task684_online_privacy_policy_text_information_type_generation
1746 task1426_country_independence_year
1747 task126_scan_structured_text_generation_command_action_all
1748 task605_find_the_longest_common_subsequence_in_two_lists
1749 task128_scan_structured_text_generation_command_action_short
1750 task960_ancora-ca-ner_named_entity_recognition
1751 task078_all_elements_except_last_i
1752 task1427_country_region_in_world
1753 task063_first_i_elements
1754 task956_leetcode_420_strong_password_check
1755 task683_online_privacy_policy_text_purpose_answer_generation
1756 task091_all_elements_from_index_i_to_j
1757 task1542_every_ith_element_from_starting
1758 task1506_celebrity_minimal_dob_span
1759 task245_check_presence_in_set_intersection
1760 task497_extract_all_numbers_from_list_in_order
1761 task1428_country_surface_area
1762 task092_check_prime_classification
1763 task1088_array_of_products
1764 task1332_check_leap_year
1765 task127_scan_long_text_generation_action_command_all
1766 task129_scan_long_text_generation_action_command_short
1767 task1322_country_government_type
1768 task1331_reverse_array
1769 task131_scan_long_text_generation_action_command_long
1770 task371_synthetic_product_of_list
1771 task1189_check_char_in_string
1772 task208_combinations_of_list
1773 task211_logic2text_classification
1774 task1551_every_ith_element_from_kth_element
1775 task1194_kth_largest_element
1776 task101_reverse_and_concatenate_all_elements_from_index_i_to_j
Rank Task
1777 task1404_date_conversion
1778 task504_count_all_alphabetical_elements_in_list
1779 task087_new_operator_addsub_arithmetic
1780 task850_synthetic_longest_palindrome
1781 task099_reverse_elements_between_index_i_and_j
1782 task206_collatz_conjecture
1783 task505_count_all_numerical_elements_in_list
1784 task1405_find_median
1785 task267_concatenate_and_reverse_all_elements_from_index_i_to_j
1786 task207_max_element_lists
1787 task1443_string_to_number
1788 task1188_count_max_freq_char
1789 task212_logic2text_classification
1790 task374_synthetic_pos_or_neg_calculation
1791 task1190_add_integer_to_list
1792 task243_count_elements_in_set_intersection
1793 task636_extract_and_sort_unique_alphabets_in_a_list
1794 task124_conala_pair_averages
1795 task1150_delete_max_min
1796 task755_find_longest_substring_and_replace_its_sorted_lowercase_version_in_both_lists
1797 task100_concatenate_all_elements_from_index_i_to_j
1798 task372_synthetic_palindrome_numbers
1799 task1148_maximum_ascii_value
1800 task506_position_of_all_alphabetical_elements_in_list
1801 task373_synthetic_round_tens_place
1802 task367_synthetic_remove_floats
1803 task1406_kth_smallest_element
1804 task1333_check_validity_date_ddmmyyyy
1805 task244_count_elements_in_set_union
1806 task205_remove_even_elements
1807 task1320_country_domain_tld
1808 task123_conala_sort_dictionary
1809 task122_conala_list_index_addition
1810 task076_splash_correcting_sql_mistake
1811 task094_conala_calculate_mean
1812 task507_position_of_all_numerical_elements_in_list
1813 task1403_check_validity_date_mmddyyyy
1814 task1315_find_range_array
1815 task098_conala_list_intersection
1816 task1087_two_number_sum
1817 task095_conala_max_absolute_value
1818 task1089_check_monotonic_array
1819 task077_splash_explanation_to_sql
1820 task097_conala_remove_duplicates
1821 task125_conala_pair_differences
1822 task368_synthetic_even_or_odd_calculation
1823 task1151_swap_max_min
1824 task852_synthetic_multiply_odds
1825 task606_sum_of_all_numbers_in_list_between_positions_i_and_j
1826 task090_equation_learner_algebra
1827 task1446_farthest_integers
1828 task096_conala_list_index_subtraction
1829 task868_cfq_mcd1_explanation_to_sql
1830 task369_synthetic_remove_odds
1831 task093_conala_normalize_lists
1832 task370_synthetic_remove_divisible_by_3
1833 task851_synthetic_multiply_evens
1834 task637_extract_and_sort_unique_digits_in_a_list
1835 task1498_24hour_to_12hour_clock
1836 task869_cfq_mcd1_sql_to_explanation
1837 task1445_closest_integers
1838 task1444_round_power_of_two
1839 task366_synthetic_return_primes
1840 task107_splash_question_to_sql
Table 7: List of 128 least representative tasks in FLAN-2022 collection as ordered by the Graph Cut
As pointed out in Section 2, the three submodular functions in Table 1 are different. Facility Location predominantly models representation; Graph Cut models a trade-off between representation and diversity; Log Determinant predominantly models diversity. To better visualize this, we present t-SNE plots of 1840 tasks present in the FLAN 2022 collection (Longpre et al., 2023) and highlight the tasks selected by these functions for different values of $M^{\prime}$. Figure 6, Figure 7 and Figure 8 respectively contains the visualizations for Facility Location, Graph Cut and Log Determinant.